# FP4 All the Way: Fully Quantized Training of LLMs

### v1

@@ 1 Introduction
Until recently, the dominant numerical format for pretraining Large Language Models (LLMs) was BF16, which provided a balance between precision and efficiency. However, as model sizes and dataset scales have grown, researchers have explored lower-precision alternatives to improve computational efficiency and reduce memory requirements. A few pioneering studies have demonstrated that full training in FP8 is not only feasible but also effective. [13] showcased the potential of FP8 training on small-scale datasets (100B),[9] extended it to trillions of tokens dataset with a 7B parameters model and [7] demonstrated FP8’s viability at an even larger scale, successfully training a massive 671B parameter Mixture-of-Experts (MoE) model on a vast dataset, achieving state-of-the-art results. FP8 is quickly emerging as the new standard for large-scale LLM training. As research continues to push precision boundaries further, the next logical step is exploring FP4, which promises even greater efficiency while maintaining training stability and model accuracy.
@@
Recently, NVIDIA introduced its new GPU architecture, Blackwell [1], marking a major milestone as the first GPU to support FP4 matrix multiplication in hardware. This advancement enables nearly 2× acceleration in matrix multiplication throughput compared to FP8, significantly boosting performance and energy efficiency for large-scale deep learning workloads. Blackwell supports two distinct FP4 formats—MXFP4 and NVFP4—each with different design choices in block size and scale encoding. In this paper, we aim to systematically investigate the full range of block sizes and scaling formats for FP4 training, analyzing their impact on accuracy and stability and highlighting the advantages of the NVFP4 format.
@@
Beyond the choice of numerical format and block size, another critical aspect of low-precision quantization is the rounding mode used during quantization. The two most commonly used rounding modes are round-to-nearest (RtN), which deterministically maps each value to its closest representable value, and stochastic rounding (SR), which probabilistically rounds based on the distance to neighboring values to reduce quantization bias. While prior empirical work [19, 5] has highlighted the importance of SR, particularly in the backward pass for stabilizing gradient updates, its benefits have largely been observed heuristically. In this work, we provide an analysis of noisy gradients training and identify the point where they stop being effective and show the importance of SR in low-precision training.
@@
Recent studies have explored the use of FP4 formats for training large language models, yet none achieved full quantization across all key components—weights, activations, and gradients. In contrast, our approach introduces, for the first time, a fully quantized FP4 training framework that addresses all these components simultaneously on significantly larger datasets. Table 2 highlights these important distinction. This end-to-end FP4 training setup enables us to evaluate the practical viability of low-precision methods at scale. Moreover, we evaluate our FP4 models on various downstream tasks—showing on-par results with the BF16 baseline—demonstrating that efficiency gains do not come at the cost of quality.
@@
    FP4 Format Optimization: We conduct comprehensive experiments on block size and scaling formats for FP4 training, revealing that block sizes below 16 elements offer diminishing returns on accuracy. Our comparison of different exponent-mantissa configurations (E1M6 through E8M0) found the best performance was achieved by E4M3—the format used in NVFP4, validating NVIDIA’s hardware design choices.
@@
    Split Rounding Strategy: We establish that combining stochastic rounding in the backward pass with round-to-nearest in the forward pass significantly improves training stability and final model accuracy in FP4 training compared to using either method alone throughout the training process.
@@
    Precision Transition Analysis: We provide a theoretical framework that identifies the critical point at which FP4 precision becomes less effective for continued training progress. Specifically, when the full-precision gradient standard deviation falls below $\sqrt{3}$ times the quantization noise standard deviation. We suggest, at the end of the training, to apply a higher precision Quantization Aware Finetuning (QAF) to increase the signal-to-noise ratio above this threshold and converge to the baseline.
@@
    End-to-End Large-Scale FP4 Training: We successfully train a 7B-parameter LLM entirely in FP4 precision using 256 Intel Gaudi2 accelerators. While a slight gap in final training loss is observed compared to the BF16 baseline, it is fully closed through a brief quantization-aware finetuning (QAF) phase. This leads to downstream task performance on par with BF16, confirming the practical viability of FP4 for real-world large-scale applications.
@@ 2 Related work
Quantization is a key area of research in the effort to compress neural networks for more efficient deployment and reduced resource consumption. The two predominant approaches in this space are Post-Training Quantization (PTQ) [22, 12, 10] and Quantization-Aware Training (QAT) [8, 4]. PTQ focuses on converting pre-trained models to low-bit representations without additional training, making it attractive for rapid deployment, especially in inference scenarios. In contrast, QAT incorporates quantization effects during model training or fine-tuning, enabling the network to adapt and maintain accuracy under low-precision constraints. Both methods have seen significant advancements, with recent work demonstrating competitive performance at 4-bit precision [10] and below [20].
@@
Fully Quantized Training (FQT) is a more challenging task than PTQ or QAT, as it requires training from scratch with low-precision weights, activations, and gradients to accelerate all matrix multiplications. Until recently, applying FQT beyond 16-bit precision was considered difficult due to instability and convergence issues. However, recent works have demonstrated its feasibility. [13] presented the first FQT of a large language model in FP8 on a dataset of up to 100 billion tokens. [9] extended this to 2 trillion tokens, revealing stability issues in later training stages and proposing a modified activation function to address them. [7] further advanced the field by training a large Mixture-of-Experts (MoE) model with FP8 FQT, mitigating instability through finer-grained quantization.
@@
Finer granularity quantization is emerging as a key direction for enabling Fully Quantized Training (FQT) beyond FP8, particularly in the context of FP4 precision. By reducing the quantization block size, these methods aim to better capture local variations in data distributions, improving stability and accuracy. Notable examples include MXFP4 [15] and NVFP4 [1].
@@
The two works most closely related to ours are [21, 19]. [21] proposes training large language models using a vector-wise FP4 format combined with two key techniques: a Differentiable Gradient Estimator (DGE) to replace the standard Straight-Through Estimator (STE), and Outlier Clamp Compensation (OCC) to handle activation outliers wih an additional sparse residual matrix. Their models are trained on up to 100 billion tokens, but they quantize only weights and activations, keeping gradients in higher precision—thus accelerating only one of the three matrix multiplications involved in training. [19], in contrast, focuses on gradient quantization using the MXFP4 format and applies stochastic rounding alongside the Hadamard transform to stabilize training. They train models up to 40 billion tokens but, like [21], only accelerate one of the three matrix multiplications. In contrast, our work is the first to demonstrate full FP4 Fully Quantized Training (FQT) of large-scale LLMs, enabling the acceleration of all matrix multiplications during training.
@@ 3 FP4 training
Going beyond FP8 to FP4 training presents significant challenges due to the limited dynamic range of FP4, making it difficult to capture the full variability of activations and gradients without excessive quantization error. However, the recently introduced microscaling floating-point family (MXFP) [15] offers a promising alternative by dynamically adjusting the scale at finer granularity, mitigating precision loss.
MXFP4, includes 1 sign bit, 2 exponent bits and 1 mantissa bit (E2M1) is a floating-point format that enhances low-precision training by dividing data into blocks of size 32, with each block sharing a common scale. The scale for each block is stored using the E8M0 format, an 8-bit exponent-only representation that provides a wide dynamic range without a mantissa and sign. Another potential format for FP4 training is NVFP4, which uses the same E2M1 data representation as MXFP4 but differs in block size and scaling format. NVFP4 divides data into smaller blocks of size 16, compared to MXFP4’s 32, allowing for finer-grained scaling adjustments. Additionally, NVFP4 employs an E4M3 format for storing scales, providing a balance between dynamic range and precision. Both MXFP4 and NVFP4 are supported in NVIDIA’s Blackwell architecture [1]. In Table 1 we compare these 2 formats.
@@ 3.1 Exploring block size and scale format
In Fig. 1 we train a 350m Llama-style model with FP4 format (E2M1) with block size 16 and different scaling formats. Note that, similar to NVFP4, most configurations (except for E8M0, which corresponds to MXFP4) do not utilize the sign bit in the scale. This may represent a potential inefficiency that future work could aim to exploit. We notice that the best results are achieved with E3M4 and E4M3, where the latter is used in the NVFP4 format. In Fig. 2 we compare different block sizes both with scales formats E8M0 and E4M3, which are the scales used in MXFP4 and NVFP4, respectively. Note that while selecting the appropriate scale has a significant impact on final accuracy—for example, E1M6 leads to complete divergence—the block size has a more modest effect, with smaller block sizes generally yielding better results. This leads us to proceed with the NVFP4 format, which uses a block size of 16 and a scale format of E4M3.
@@ 3.2 Exploring the rounding modes
[…]
where $C$ is the loss function, $\odot$ is a component-wise product and, in each layer $l$, $f_{l}$ is the activation function, $W_{l}$ is weight matrix, $z_{l}$ are the pre-activations, and $g_{l}\triangleq\frac{\partial C}{\partial a_{l}}$.
@@
Notably, we have the flexibility to select the rounding mode independently for each of these six elements. In Fig. 3, we present results of training a 350m Llama model using NVFP4 where stochastic rounding (SR) is applied to each of these elements separately while the rest of the elements use round-to-nearest (RtN), allowing us to evaluate its individual contribution. Applying SR to neural gradients during update and backward GEMMs and activations during the update GEMM helps reduce training loss, whereas using SR in other components leads to an increase in loss. As a result, we adopt the following selective rounding scheme:
[…]
@@ 4 Analysis of Quantized SGD with Stochastic Rounding
  • SR enables unbiased updates, allowing stable descent even under aggressive FP4 quantization. In contrast, deterministic rounding introduces a persistent bias, leading to an irreducible error floor and preventing convergence (see Section B.2).
  • There exists a critical threshold: With SR, we show in Section 4.1 (and in more detail in Section B.1) that the average per-coordinate gradient magnitude falls approximately below $\sqrt{3}$ times the quantization noise standard deviation, training no longer yields effective loss reduction. This threshold is derived under simplifying assumptions—e.g., gradient descent with optimal step size, Taylor approximation of the loss, and a concentrated Hessian spectrum.
  • Empirical evidence supports the theory: in Section 4.2, for both synthetic and real-model settings, we show empirically performance degrades sharply below this threshold. A precision switch guided by the theory restores convergence.
@@ 4.1 Theoretical Derivation
We begin with a second-order Taylor expansion of the loss function around the current parameter vector $\theta_{t}$, which reveals a descent term proportional to $-\nabla L^{T}\Delta\theta$ and a curvature term involving the Hessian $H$. We then replace the full-precision gradient $\nabla L$ with its quantized version $g_{q}=\nabla L+\varepsilon$, where $\varepsilon$ is zero-mean noise introduced by stochastic rounding.
@@ 4.2 Empirical Validation
First, we simulate training on a simple quadratic loss with an adaptive noise schedule. We scale the quantization noise to $\sigma_{q}=k\cdot\sigma_{\mathrm{critical}}$ for $k=2,1,0.5$. As shown in Fig. 4, convergence completely stalls at high noise levels (e.g., $k=2$), slows near the critical threshold ($k=1$), and closely tracks full-precision training when noise is reduced below the threshold ($k=0.5$).
@@
Second, we test in Fig. 5, the threshold on a real 60M-parameter Llama model. During training, we monitor the ratio $\|\nabla L\|/(\sigma_{q}\sqrt{d})$, and switch to higher-precision gradients at the 1000th iteration. When this ratio crosses $\sqrt{3}$, the loss gap to the full-precision baseline closes immediately, validating the predictive power of the threshold.
@@ 5 Experiments
We used the Llama2 model [18] as our baseline. This model is a decoder-only Transformer [3] with pre-normalization RMSNorm [23], Smooth-SwiGLU activation function [9], and rotary positional embeddings [17]. We trained the models on the open-source Red Pajama dataset [6] for $\sim 200B$ tokens, maintaining hyperparameters consistent with [18], including train-test split and initialization. All training was conducted on 256 Intel Gaudi2 devices, during $\sim 6$ days. The training is still in progress; we will upload a new version using the larger dataset once it is complete.
@@
In Fig. 6(a) we present our main experiment and present the training loss of Llama2 7B with the proposed FP4 scheme, which includes the use of the NVFP4 format (block size 16, scale format E4M3) and applying SR in the neural gradients (update + backward GEMMs) and activations (update GEMM), while applying RtN for the weights (forward + backward GEMMs) and activations (forward GEMM). We observe a slight difference in training loss that does not translate into a meaningful gap in downstream task performance (Table 3), consistent with findings reported in prior work [2]. In Table 2 we compare the quantization settings of our work with two previous FP4 training works [19, 21], showing we are the first work that allows the acceleration of all matrix multiplication during training. In appendix LABEL:fig:fullFp4App we continue the training beyond the point reached at the submission deadline.
Figure 6: Full FP4 trainng a 7B model achieved same loss as BF16 baseline. (a): Training loss of Llama2 7B using the proposed FP4 scheme which include NVFP4 format (block size 16, scale format E4M3) with SR in the neural gradients (update + backward GEMMs) and activations (update GEMM). Notice a small gap in training loss is observed, but it does not result in a significant difference in downstream task performance (Table 3). (b): Quanization aware finetuning (QAF) training loss of Llama2 7B using NVFP4 format in the forward GEMM and BF16 in the backward and update GEMMs. Notice this short QAF is able to close completely the gap with BF16 baseline.
Table 2: Comparison of the different FP4 training works. DGE refers to Differentiable Gradient Estimator, OCP to Outlier Clamp Compensation, SR to Stochastic Rounding and RHT to Random Hadamard Transform. While previous works quantize only part of the GEMMs to FP4 with additional overhead of residual sparse matrix (OCC [21]) or Hadamard transform (RHT [19]), our work is the first work that show quantization to FP4 of all GEMMs operands, without adding significant overhead.
@@
[21] FP4+DGE+OCC FP4+DGE+OCC BF16 100B
[19] BF16 BF16 MXFP4+RHT+SR 21B
Ours NVFP4 (RtN) NVFP4 (RtN / SR) NVFP4 (SR) 200B
@@
The FP4 training results (Fig. 6(a)) in a small gap in training loss compared to the BF16 baseline. As shown in Fig. 5, increasing the precision raises the gradient-to-noise ratio above the critical threshold. To close the remaining gap, we introduce a brief quantization-aware finetuning (QAF) phase, where pretraining continues on the same dataset with the forward pass kept in FP4, while the backward pass is executed in BF16. Keeping the forward path in FP4 ensures that the model remains fully compatible with low-precision inference. During this stage, we reset the learning rate and apply a short warmup (40 iterations) followed by a cosine decay schedule with an initial peak learning rate. In Fig. 6(b) we present that this short QAF can close completely the gap with BF16 baseline.
@@
Table 3 compares the zero-shot performance (accuracy and perplexity) on downstream tasks between the BF16 baseline and our FP4 model — both after the FP4 training phase (200B tokens) and after QAF phase (+20B tokens). Notice that while a small gap is observed in part of the tasks after the FP4 training phase, it is completely closed after the QAF.
Table 3: Zero shot accuracy and perplexity comparison between the BF16 baseline and the proposed FP4, both after the FP4 training phase (200B) and after QAF phase (+20B). HS refers to HellaSwag, WG refers to Winogrande, BQ refers to BoolQ, AC refers to Arc-C, PQ refers to PiQA, LA refers to Lambada. Notice that after the QAF the proposed FP4 model achieved on-par results with the BF16 baseline.
Precision Tokens Accuracy $\uparrow$ Perplexity $\downarrow$
LA HS WG AC BQ PQ Avg Wiki LA
BF16 200B 54.51 62.12 60.69 32.94 63.27 73.88 57.9 6.29 8.48
BF16 220B 53.81 62.1 62.12 33.11 63.58 73.34 58.01 6.21 8.68
FP4 200B 54.45 60.78 61.56 33.96 60.12 73.56 57.41 6.5 8.46
+ QAF +20B 55.44 62.29 61.72 34.56 64.31 74.37 58.78 6.24 7.62

### v2

@@ 1 Introduction
Until recently, the dominant numerical format for pretraining Large Language Models (LLMs) was BF16, which provided a balance between precision and efficiency. However, as model sizes and dataset scales have grown, researchers have explored lower-precision alternatives to improve computational efficiency and reduce memory requirements. A few pioneering studies have demonstrated that full training in FP8 is not only feasible but also effective. (Peng et al., 2023) showcased the potential of FP8 training on small-scale datasets (100B),Fishman et al. (2024) extended it to trillions of tokens dataset with a 7B parameters model, and (DeepSeek-AI et al., 2024) demonstrated FP8’s viability at an even larger scale, successfully training a massive 671B parameter Mixture-of-Experts (MoE) model on a vast dataset, achieving state-of-the-art results. FP8 is quickly emerging as the new standard for large-scale LLM training. As research continues to push precision boundaries further, the next logical step is exploring FP4, which promises even greater efficiency while maintaining training stability and model accuracy.
@@
Recently, NVIDIA introduced its new GPU architecture, Blackwell (Nvi, ), marking a major milestone as the first GPU to support FP4 matrix multiplication in hardware. This advancement enables nearly 2× acceleration in matrix multiplication throughput compared to FP8, significantly boosting performance and energy efficiency for large-scale deep learning workloads. Blackwell supports two distinct FP4 formats—MXFP4 and NVFP4—each with different design choices in block size and scale encoding. In this paper, we aim to systematically investigate the full range of block sizes and scaling formats for FP4 training, analyzing their impact on accuracy and stability and highlighting the advantages of the NVFP4 format.
@@
Beyond the choice of numerical format and block size, another critical aspect of low-precision quantization is the rounding mode used during quantization. The two most commonly used rounding modes are round-to-nearest (RtN), which deterministically maps each value to its closest representable value, and stochastic rounding (SR), which probabilistically rounds based on the distance to neighboring values to reduce quantization bias. While prior empirical work Tseng et al. (2025); Chmiel et al. (2021) has highlighted the importance of SR, particularly in the backward pass for stabilizing gradient updates, its benefits have largely been observed heuristically. In this work, we provide an analysis of noisy gradients training and identify the point where they stop being effective and show the importance of SR in low-precision training.
@@
Recent studies have explored the use of FP4 formats for training large language models, yet none achieved full quantization across all key components—weights, activations, and gradients. In contrast, our approach introduces, for the first time, a fully quantized FP4 training framework that addresses all these components simultaneously, and on significantly larger datasets. Table˜2 highlights these important distinction. This end-to-end FP4 training setup enables us to evaluate the practical viability of low-precision methods at scale. Moreover, we evaluate our FP4 models on various downstream tasks—showing on-par results with the BF16 baseline—demonstrating that efficiency gains do not come at the cost of quality.
@@
    FP4 Format Optimization. We conduct comprehensive experiments on block size and scaling formats for FP4 training, revealing that block sizes below 16 elements offer diminishing returns on accuracy. Our comparison of different exponent-mantissa configurations (E1M6 through E8M0) found the best performance was achieved by E4M3—the format used in NVFP4, validating NVIDIA’s hardware design choices.
@@
    Split Rounding Strategy. We establish that combining stochastic rounding in the backward pass with round-to-nearest in the forward pass significantly improves training stability and final model accuracy in FP4 training compared to using either method alone throughout the training process.
@@
    Precision Transition Analysis. We provide a theoretical framework that identifies the critical point at which FP4 precision becomes less effective for continued training progress. Specifically, when the full-precision gradient standard deviation falls below $\sqrt{3}$ times the quantization noise standard deviation. We suggest, at the end of the training, to apply a higher precision Quantization Aware Finetuning (QAF) to increase the signal-to-noise ratio higher above this threshold and quickly converge to the baseline.
@@
    End-to-End Large-Scale FP4 Training. We successfully train a 7B-parameter LLM entirely in FP4 precision on a trillion tokens, using 256 Intel Gaudi2 accelerators. While a slight gap in final training loss is observed compared to the BF16 baseline, it is fully closed through the brief QAF phase. This leads to downstream task performance on par with BF16, confirming the practical viability of FP4 for real-world large-scale applications.
@@ 2 Related work
Quantization is a key area of research in the effort to compress neural networks for more efficient deployment and reduced resource consumption. The two predominant approaches in this space are Post-Training Quantization (PTQ) (Xiao et al., 2022; Lin et al., 2023; Frantar et al., 2022) and Quantization-Aware Training (QAT) (Dettmers et al., 2023; Cheng et al., 2023; Panferov et al., 2025). PTQ focuses on converting pre-trained models to low-bit representations without additional training, making it attractive for rapid deployment, especially in inference scenarios. In contrast, QAT incorporates quantization effects during model training or fine-tuning, enabling the network to adapt and maintain accuracy under low-precision constraints. Both methods have seen significant advancements, with recent work demonstrating competitive performance at 4-bit precision Frantar et al. (2022) and below Wang et al. (2023).
@@
Fully Quantized Training (FQT) is a more challenging task than PTQ or QAT, as it requires training from scratch with low-precision weights, activations, and gradients to accelerate all matrix multiplications. Until recently, applying FQT beyond 16-bit precision was considered difficult due to instability and convergence issues. However, recent works have demonstrated its feasibility. Peng et al. (2023) presented the first FQT of a large language model in FP8 on a dataset of up to 100 billion tokens. Fishman et al. (2024) extended this to 2 trillion tokens, revealing stability issues in later training stages and proposing a modified activation function to address them. (DeepSeek-AI et al., 2024) further advanced the field by training a large Mixture-of-Experts (MoE) model with FP8 FQT, mitigating instability through finer-grained quantization.
@@
Finer granularity quantization is emerging as a key direction for enabling Fully Quantized Training (FQT) beyond FP8, particularly in the context of FP4 precision. By reducing the quantization block size, these methods aim to better capture local variations in data distributions, improving stability and accuracy. Notable examples include MXFP4 Rouhani et al. (2023) and NVFP4 Nvi .
@@
The two works most closely related to ours are Wang et al. (2025); Tseng et al. (2025). Wang et al. (2025) proposes training large language models using a vector-wise FP4 format combined with two key techniques: a Differentiable Gradient Estimator (DGE) to replace the standard Straight-Through Estimator (STE), and Outlier Clamp Compensation (OCC) to handle activation outliers wih an additional sparse residual matrix. Their models are trained on up to 100 billion tokens, but they quantize only weights and activations, keeping gradients in higher precision—thus accelerating only one of the three matrix multiplications involved in training. Tseng et al. (2025), in contrast, focuses on gradient quantization using the MXFP4 format and applies stochastic rounding alongside the Hadamard transform to stabilize training. They train models up to 40 billion tokens. However, like Wang et al. (2025), they only accelerate part of the three matrix multiplications. In contrast, our work is the first to demonstrate full FP4 Fully Quantized Training (FQT) of large-scale LLMs, enabling the acceleration of all matrix multiplications during training.
@@ 3 FP4 training
Going beyond FP8 to FP4 training presents significant challenges due to the limited dynamic range of FP4, making it difficult to capture the full variability of activations and gradients without excessive quantization error. However, the recently introduced microscaling floating-point family (MXFP) Rouhani et al. (2023) offers a promising alternative by dynamically adjusting the scale at finer granularity, mitigating precision loss.
MXFP4, includes 1 sign bit, 2 exponent bits, and 1 mantissa bit (E2M1) is a floating-point format that enhances low-precision training by dividing data into blocks of size 32, with each block sharing a common scale. The scale for each block is stored using the E8M0 format, an 8-bit exponent-only representation that provides a wide dynamic range without a mantissa and sign. Another potential format for FP4 training is NVFP4, which uses the same E2M1 data representation as MXFP4 but differs in block size and scaling format. NVFP4 divides data into smaller blocks of size 16, compared to MXFP4’s 32, allowing for finer-grained scaling adjustments. Additionally, NVFP4 employs an E4M3 format for storing scales, providing a balance between dynamic range and precision. Both MXFP4 and NVFP4 are supported in NVIDIA’s Blackwell architecture Nvi . In Table˜1 we compare these 2 formats.
@@ 3.1 Exploring block size and scale format
In Fig.˜1 we train a 350M Llama-style model with FP4 format (E2M1) with block size 16 and different scaling formats. Note that, similar to NVFP4, most configurations (except for E8M0, which corresponds to MXFP4) do not utilize the sign bit in the scale. This may represent a potential inefficiency that future work could aim to exploit. We notice that the best results are achieved with E3M4 and E4M3, where the latter is used in the NVFP4 format. In Fig.˜2 we compare different block sizes both with scales formats E8M0 and E4M3, which are the scales used in MXFP4 and NVFP4, respectively. Note that while selecting the appropriate scale has a significant impact on final accuracy—for example, E1M6 leads to complete divergence—the block size has a more modest effect, with smaller block sizes generally yielding better results. This leads us to proceed with the NVFP4 format, which uses a block size of 16 and a scale format of E4M3.
@@ 3.2 Exploring the rounding modes
$\textbf{[Forward]}\quad z_{l}=Q(W_{l})Q(a_{l-1});\quad\quad a_{l}=f_{l}(z_{l})$ (1)
$\textbf{[Backward]}\quad g_{l-1}=Q(W_{l}^{T})Q(\delta_{l});\quad\delta_{l}=f_{l}^{\prime}(z_{l})\odot g_{l}$ (2)
$\textbf{[Update]}\quad\frac{\partial C}{\partial W_{l}}=Q(\delta_{l})Q(a_{l-1}^{T})\,$, (3)
where $C$ is the loss function, $Q$ is a quantization operation, $\odot$ is a component-wise product and, in each layer $l$, $f_{l}$ is the activation function, $W_{l}$ is weight matrix, $z_{l}$ are the pre-activations, and $g_{l}\triangleq\frac{\partial C}{\partial a_{l}}$.
@@
Notably, we have the flexibility to select the rounding mode independently for each of these six elements. In Fig.˜3, we present results of training a 350M Llama model using NVFP4 where stochastic rounding (SR) is applied to each of these elements separately while the rest of the elements use round-to-nearest (RtN), allowing us to evaluate its individual contribution. Applying SR to neural gradients during update and backward GEMMs and activations during the update GEMM helps reduce training loss, whereas using SR in other components leads to an increase in loss. In Fig.˜7 in the Appendix, we present additional experiments that support the same conclusion. As a result, we adopt the following selective rounding scheme:
$\textbf{[Forward]}\quad z_{l}=Q_{\mathrm{RtN}}(W_{l})Q_{\mathrm{RtN}}(a_{l-1});\quad\quad a_{l}=f_{l}(z_{l})$ (4)
$\textbf{[Backward]}\quad g_{l-1}=Q_{\mathrm{RtN}}(W_{l}^{T})Q_{\mathrm{SR}}(\delta_{l});\quad\delta_{l}=f_{l}^{\prime}(z_{l})\odot g_{l}$ (5)
$\textbf{[Update]}\quad\frac{\partial C}{\partial W_{l}}=Q_{\mathrm{SR}}(\delta_{l})Q_{\mathrm{SR}}(a_{l-1}^{T})\,$, (6)
@@ 4 Analysis of Quantized SGD with Stochastic Rounding
  • SR enables unbiased updates, allowing stable descent even under aggressive FP4 quantization. In contrast, deterministic rounding introduces a persistent bias, leading to an irreducible error floor and preventing convergence (see Section˜B.2).
  • There exists a critical threshold: With SR, we show in Section˜4.1 (and in more detail in Section˜B.1) that the average per-coordinate gradient magnitude falls approximately below $\sqrt{3}$ times the quantization noise standard deviation, training no longer yields effective loss reduction. This threshold is derived under simplifying assumptions—e.g., gradient descent with optimal step size, Taylor approximation of the loss, and a concentrated Hessian spectrum.
  • Empirical evidence supports the theory: in Section˜4.2, for both synthetic and real-model settings, we show empirically performance degrades sharply below this threshold. A precision switch guided by the theory restores convergence.
@@ 4.1 Theoretical Derivation
We begin with a second-order Taylor expansion of the loss function around the current parameter vector $\theta_{t}$, which reveals a descent term proportional to $\nabla L^{T}\Delta\theta$ and a curvature term involving the Hessian $H$. We then replace the full-precision gradient $\nabla L$ with its quantized version $g_{q}=\nabla L+\varepsilon$, where $\varepsilon$ is zero-mean noise introduced by stochastic rounding.
@@ 4.2 Empirical Validation
First, we simulate training on a simple quadratic loss with an adaptive noise schedule. We scale the quantization noise to $\sigma_{q}=k\cdot\sigma_{\mathrm{critical}}$ for $k=2,1,0.5$. As shown in Fig.˜4, convergence completely stalls at high noise levels (e.g., $k=2$), slows near the critical threshold ($k=1$), and closely tracks full-precision training when noise is reduced below the threshold ($k=0.5$).
@@
Second, we test in Fig.˜5, the threshold on a real 60M-parameter Llama model. During training, we monitor the ratio $\|\nabla L\|/(\sigma_{q}\sqrt{d})$, and switch to higher-precision gradients at the 1000th iteration. When this ratio crosses $\sqrt{3}$, the loss gap to the full-precision baseline closes immediately, validating the predictive power of the threshold.
@@ 5 Experiments
We used the Llama2 model (Touvron et al., 2023) as our baseline. This model is a decoder-only Transformer (Brown et al., 2020) with pre-normalization RMSNorm (Zhang and Sennrich, 2019), Smooth-SwiGLU activation function (Fishman et al., 2024), and rotary positional embeddings (Su et al., 2024). We trained the models on the open-source Red Pajama dataset (Computer, 2023) for $1T$ tokens, maintaining hyperparameters consistent with Touvron et al. (2023), including train-test split and initialization. Specifically, we used AdamW optimizer with $\beta_{1}=0.9$, $\beta_{2}=0.95$. We used cosine learning rate schedule, with 2000 steps of warmup, peak learning rate of $3\times 10^{-4}$ and decay to 0.1 of the peak learning rate. We used a global batch-size of 4M tokens. All training was conducted on 256 Intel Gaudi2 devices, during $\sim 30$ days.
@@
In Fig.˜6(a) we present our main experiment and present the training loss of Llama2 7B with the proposed FP4 scheme, which includes the use of the NVFP4 format (block size 16, scale format E4M3) and applying SR in the neural gradients (update + backward GEMMs) and activations (update GEMM), while applying RtN for the weights (forward + backward GEMMs) and activations (forward GEMM). In Table˜2 we compare the quantization settings of our work with two previous FP4 training works Tseng et al. (2025); Wang et al. (2025), showing we are the first work that allows the acceleration of all matrix multiplication during training. In the Appendix Table˜4 we show the similarity of the FP4 training losses at different seeds, showing the noise robustness of the proposed method.
Figure 6: Full FP4 training a 7B model achieved same loss as BF16 baseline. (a): Training loss of Llama2 7B using the proposed FP4 scheme which include NVFP4 format (block size 16, scale format E4M3) with SR in the neural gradients (update + backward GEMMs) and activations (update GEMM). Notice a small gap in training loss is observed, but it does not result in a significant difference in downstream task performance (Table˜3). (b): Quanization aware finetuning (QAF) training loss of Llama2 7B using NVFP4 format in the forward GEMM and BF16 in the backward and update GEMMs. Notice this short QAF is able to completely close the gap with BF16 baseline.
Table 2: Comparison of the different FP4 training works. DGE refers to Differentiable Gradient Estimator, OCC to Outlier Clamp Compensation, SR to Stochastic Rounding and RHT to Random Hadamard Transform. While previous works quantize only part of the GEMMs to FP4 with additional overhead of residual sparse matrix (OCC Wang et al. (2025)) or Hadamard transform (RHT Tseng et al. (2025)), our work is the first work that show quantization to FP4 of all GEMMs operands, without adding significant overhead.
@@
Wang et al. (2025) FP4+DGE+OCC FP4+DGE+OCC BF16 100B
Tseng et al. (2025) BF16 BF16 MXFP4+RHT+SR 21B
Ours NVFP4 (RtN) NVFP4 (RtN / SR) NVFP4 (SR) 1T
@@
FP4 training results (Fig.˜6(a)) in a small gap in training loss compared to the BF16 baseline. As shown in Fig.˜5, increasing the precision raises the gradient-to-noise ratio higher above the critical threshold. To close the remaining gap, we introduce a brief quantization-aware finetuning (QAF) phase, where pretraining continues on the same dataset with the forward pass kept in FP4, while the backward pass is executed in BF16. Keeping the forward path in FP4 ensures that the model remains fully compatible with low-precision inference, without requiring any additional processes such as post-training-quantization. During this stage, we reset the learning rate and apply a short warmup (40 iterations) followed by a cosine decay schedule with an initial peak learning rate. In Fig.˜6(b) we present that this short QAF can completely close the gap with BF16 baseline, achieving an average bits of 4.3 bits for GEMMs across training + QAF. In the Appendix Table˜5 we show an ablation study of the QAF length ratio required to get similar final loss as BF16. We find it decreases for larger datasets.
@@
Table 3 compares the zero-shot performance (accuracy and perplexity) on downstream tasks between the BF16 baseline and our FP4 model — both after the FP4 training phase (1T tokens) and after QAF phase (+40B tokens). Notice that while a small gap is observed in part of the tasks after the FP4 training phase, it is completely closed after the QAF.
Table 3: Zero shot accuracy and perplexity comparison between the BF16 baseline and the proposed FP4, both after the FP4 training phase (1T) and after QAF phase (+40B). HS refers to HellaSwag, WG refers to Winogrande, BQ refers to BoolQ, AC refers to Arc-C, PQ refers to PiQA, LA refers to Lambada, GQ refers to GPQA. IE refers to If Eval. MB refers to MBPP, TQ refers to TrivialQA. XS refers to XSum. Notice that after the QAF, the proposed FP4 model achieved on-par results with the BF16 baseline.
Precision Data Accuracy $\uparrow$ Perplexity $\downarrow$
LA HS WG AC BQ PQ GQ IE MB TQ XS Avg Wiki LA
BF16 1T 61.52 68.71 66.54 38.14 69.33 76.33 24.54 33.81 8.2 34.99 11.97 45.63 5.54 6.1
BF16 1.04T 61.46 68.57 64.48 38.91 70.09 75.41 24.73 30.94 8.6 34.19 12.33 45.13 5.54 6
FP4 1T 58.39 67.31 64.01 38.65 69.33 74.86 24.73 32.13 6.4 34.3 12.09 44.46 5.83 6.77
+ QAF +40B 67.71 68.53 65.98 39.25 68.9 76.01 27.29 32.25 9.4 38.31 12.11 45.75 5.56 5.97